# UMArm: Untethered, Modular, Wearable, Soft Pneumatic Arm

### v1

@@ V-A Safe and Dexterous Arm
Having many independently controlled actuators also grants the UMArm the ability to perform delicate tasks. It can move very small distances with great precision. We showcase that ability through an object-stacking experiment. Figure 8A shows the end-effector we have developed for this task. This end-effector consists of two encoder motors, one gripper driven by a servo, and the controller circuits. The two encoder motors can maintain the gripper’s orientation. Figure 8B shows the object stacking experiment. There are three objects on the upper level of a two-level platform. The UMArm picks up and places the first object onto the lower platform, stacks the second object on top of the first object, and stacks the third object on top of both objects. In the supplementary video, the robot can keep repeating this task nonstop with great consistency.
V-B Wearable Assistive Arm
@@
UMArm is very suitable to be a wearable assistive arm because of its lightweight design, untethered operation ability, large payload capacity, and dexterity. Fig. 9A shows a worker wearing the UMArm as a worksite assistive arm. A rig built for a camera stabilizer is repurposed to attach the arm to the worker. The task for the worker is to perpendicularly join two wooden planks using two wood screws. As the worker needs to use one hand to hold the plank and another hand to hold the impact driver, thus an extra hand is needed to hold the wood screw. Figure 9 B&C demonstrates how the UMArm could assist in such a situation. The arm can use the gripper to pick up the first screw from the screw holder, and align the screw at where the worker wants it. As the worker starts to drill in the screw, the arm can automatically pick up the second screw, and align it again. With the help of UMArm, the worker is able to complete the task within one minute.
V-C Inchworm Robot
@@
The regulator-embedded design, untethered operation ability, and large number of independently controlled actuators have made the UMArm extremely versatile. We want to demonstrate the versatility by showing that the UMArm can not only serve as a robotic arm, but also operate as a standalone robot that can untetheredly locomote and interact. With some slight modifications, we turned the UMArm into an untethered inchworm robot. Figure 10A shows the modified UMArm. Four legs are installed on the UMArm. The legs are metal spikes that have a 45° angle with the ground. It has less traction sliding forward and larger traction sliding backward. The compressed air tank and 12V battery pack are loaded on a wheeled trailer, which will be tagged along with the inchworm robot. Figure 10B shows the gaits of the inchworm. It raises and steps forward the legs one at a time. When inchworming forward, it can complete one cycle in 1.7 seconds, during which it moves forward by 93 millimeters. A more comprehensive demonstration of its ability is shown in Figure 10C. The inchworm is tasked with turning on the flip switch of a lightbox. It rotates and moves toward the switch. After it aligns its head with the switch, it needs to raise the first segment to flip the switch. However, if the other two segments are relaxed, as the first segment tries to exert force at the switch, the leg between the second and third segments will be lifted up, therefore, not enough force will be exerted onto the switch. Fortunately, the robot can employ the antagonistic stiffening feature to increase the stiffness of the body. By stiffening the body, the two segments at the back will stay straight and provide firm support for the raised first segment, which can now apply enough force to turn on the switch.
@@ VI Discussion
This work demonstrates the hardware novelty of UMArm. We showed that the arm can already accomplish real-world tasks with a basic kinematic model and a PID controller. We anticipate that advanced control techniques tailored to our hardware system could better exploit the UMArm’s capabilities and provide faster and more efficient responses with smartly tuned antagonistic pressure. The arm’s dynamics and advanced controls will be explored in future work. This study also did not fully demonstrate the UMArm’s directional compliance control capabilities. The large number of compliance-controlled DoFs of the arm is designed to quantitatively allow variation in compliance across different directions, enhancing the robot’s ability to safely interact with its environment while retaining necessary support and precision. However, our experiments thus far have only qualitatively demonstrated one compliance profile. Exploring control strategies to realize and fine-tune directional compliance remains an important goal for our next work in progress. Apart from control strategies, several hardware upgrades have been identified to further enhance the UMArm’s performance and robustness. For instance, replacing the present pneumatic valves with higher pressure-rated, high-flow alternatives would enable faster actuation and greater force output. Also, increasing the number of spine segments would expand the arm’s reachable workspace and dexterity.
@@ VII Conclusion
This work introduces the UMArm, a pneumatically actuated rigid-soft hybrid arm that overcomes key limitations in soft robotic arms by integrating 24 independently controlled valve-embedded McKibben actuators. Our design achieves high precision (end-effector steady-state RMSE 2.6 mm), strong force output (3 kg payload), and rapid actuation while remaining lightweight (1.15 kg). Unlike traditional pneumatic arms requiring extensive external hardware, UMArm operates untethered using a compact, onboard pneumatic system, significantly enhancing portability. The modular architecture enables scalability and reconfiguration. The UMArm also demonstrates versatility. It can transform into an untethered inchworm robot that can move and interact. By leveraging antagonistic stiffening, the arm achieves directional compliance control. It can be tuned soft in one direction for safe human interaction, and rigid in the other direction for better force transmission. These advancements position UMArm as a robust solution for precise, wearable, and multimodal robotic applications. In the future, we would like to explore various control strategies that exploit the hardware advantages of the UMArm, striving for the goal of deploying high-performance, structurally compliant arms into the unstructured world.

### v2

@@ V-B Precision Handling Arm
Having many independently controlled actuators also grants the UMArm the ability to perform delicate tasks. It can move very small distances with great precision. We showcase that ability through an object-stacking experiment. The top-right corner of Figure 9 shows the end-effector we have developed for this task. This end-effector consists of two encoder motors, one gripper driven by a servo, and the controller circuits. The two encoder motors can maintain the gripper’s orientation. Figure 9 shows the object stacking experiment. There are three objects on the upper level of a two-level platform. The UMArm picks up and places the first object onto the lower platform, stacks the second object on top of the first object, and stacks the third object on top of both objects. In the supplementary video, the robot can keep repeating this task nonstop with great consistency.
V-C Wearable Assistive Arm
@@
UMArm is very suitable to be a wearable assistive arm because of its lightweight design, untethered operation ability, large payload capacity, and dexterity. Many previous works have explored the use of soft robots for wearable assistive arms [Nguyen2019Fabric, Nguyen2019Soft, Hussain2016SixthFinger]. However, few designs aim to achieve portability, precision, payload capacity, dexterity, and a wide range of motion simultaneously, which are the goals addressed by UMArm. Fig. 10A shows a worker wearing the UMArm as a worksite assistive arm. A rig built for a camera stabilizer is repurposed to attach the arm to the worker. The task for the worker is to perpendicularly join two wooden planks using two wood screws. As the worker needs to use one hand to hold the plank and another hand to hold the impact driver, thus an extra hand is needed to hold the wood screw. Figure 10 B&C demonstrates how the UMArm could assist in such a situation. The arm can use the gripper to pick up the first screw from the screw holder, and align the screw at where the worker wants it. As the worker starts to drill in the screw, the arm can automatically pick up the second screw, and align it again. With the help of UMArm, the worker is able to complete the task within one minute.
V-D Inchworm Robot
The regulator-embedded design, untethered operation ability, and large number of independently controlled actuators have made the UMArm extremely versatile. We want to demonstrate the versatility by showing that the UMArm can not only serve as a robotic arm, but also operate as a standalone robot that can untetheredly locomote and interact. While many soft robots have demonstrated impressive locomotion capabilities through diverse mechanisms [ahmad2018kirigami, qi2023bioinspired, qin2018design], UMArm demonstrates that untethered locomotion and interaction can be achieved simultaneously using the same set of actuators.
@@ V-D Inchworm Robot
With some slight modifications, we turned the UMArm into an untethered inchworm robot. Figure 11A shows the modified UMArm. Four legs are installed on the UMArm. The legs are metal spikes that have a 45° angle with the ground. It has less traction sliding forward and larger traction sliding backward. The compressed air tank and 12V battery pack are loaded on a wheeled trailer, which will be tagged along with the inchworm robot. Figure 11B shows the gaits of the inchworm. It raises and steps forward the legs one at a time. When inchworming forward, it can complete one cycle in 1.7 seconds, during which it moves forward by 93 millimeters. A more comprehensive demonstration of its ability is shown in Figure 11C. The inchworm is tasked with turning on the flip switch of a lightbox. It rotates and moves toward the switch. After it aligns its head with the switch, it needs to raise the first segment to flip the switch. However, if the other two segments are relaxed, as the first segment tries to exert force at the switch, the leg between the second and third segments will be lifted up, therefore, not enough force will be exerted onto the switch. Fortunately, the robot can employ the antagonistic stiffening feature to increase the stiffness of the body. By stiffening the body, the two segments at the back will stay straight and provide firm support for the raised first segment, which can now apply enough force to turn on the switch.
@@ VI Discussion
VI-A Discussion on Experimental Results
In the inverse kinematics control experiment, we evaluated the arm’s performance in moving from an initial position to a target setpoint. We observed a clear pattern: a heavier payload slows down convergence and introduces more instabilities, but increasing the structural stiffness improves settling time and enhances stability. In the force transmission experiment, the arm produced the expected results. The maximum force was comparable when the arm was straight. However, when the arm was bent (positions 2 and 3), it generated a greater pulling force than pushing force. This behavior arises from the mechanics of the McKibben actuator: when stretched, the actuator produces significantly higher pulling force. In bent configurations, more actuators operate at their optimal length for pulling, but at a less favorable length for pushing. The maximum measured force was 53.3 N, while the arm itself weighs only 1.15 kg.
In the pattern drawing experiment, we demonstrated that the robot can accurately trace prescribed paths both with and without a payload. The controller used is a PID controller, which responds to the instantaneous error between the current and target positions, rather than planning motion in advance as in model predictive control. As a result, some tracking error is inevitable. Compared with [bruder2021koopman], the UMArm achieved an average error of 19 mm when tracing a circular trajectory under no load, while the arm by Bruder et al. reported 26 mm error. Moreover, UMArm operated on a substantially larger scale (trajectory radius 250 mm vs. 100 mm), making its relative performance even stronger. The reported error represents the distance between the target and end-effector positions at the same timestamp; thus, it does not fully capture trajectory similarity, which can be better appreciated by directly inspecting the plotted trajectories. Larger errors occur at sharp corners because the target position advances at constant speed along the trajectory, while the robot’s end-effector inevitably lags. Since the trajectory is not adjusted to wait for the end-effector, the robot’s motion appears less precise at these turns.
In the variable stiffness experiment, Figure B compares the estimated displacement with the actual displacement. While the results are qualitatively similar, it is worth pointing out that they should not coincide in general. The estimated displacement corresponds to a virtual displacement under the action of a virtual force and is valid only locally. By contrast, the actual displacement vector is computed from the initial and displaced end-effector positions, which is a non-local measurement from the motion capture system. Discrepancies are therefore expected. More accurate estimation of the true displacement may be achieved with improved modeling, which may be addressed in future work.
VI-B Limitations and Future Works
Several limitations were identified during the development of UMArm. First, its workspace is relatively restricted compared to conventional rigid robot arms. Although UMArm has 12 independently actuated revolute joints, each has a limited range of motion ($\pm 0.3$ rad), resulting in a workspace resembling a shallow eggshell (workspace plot available in the Supplementary Materials). To address this limitation, a linear rail with a stepper motor could be integrated at the base to provide motion along the +z+z axis. This simple addition would significantly enlarge the workspace without diminishing key features such as variable structural stiffness.
Second, we recognize the limitations of the current PID controller. In this work, our primary goal was to demonstrate the hardware capabilities of UMArm under a simple controller. Advanced control strategies tailored to the UMArm hardware may better leverage its unique properties, achieving faster, more efficient responses by intelligently tuning the large amount of independently tunable actuator pressures.
Third, the absence of internal joint angle sensing restricts the arm to environments equipped with external motion capture. Incorporating embedded joint angle feedback would allow operation beyond the motion capture area. Specifically, we can build our robot joints using Metal 3D printing, and Hall-effect angle encoders will be integrated onto the joints. The encoder can detect the rotation angle of the magnetic shaft, thereby acquiring the joint angles.
Fourth, while UMArm currently relies on a rigid spine with revolute joints to ensure stability and precise force transmission, this architecture necessarily limits the range of continuous deformations compared to fully soft or continuum arms. It could be helpful to investigate hybrid strategies that increase structural compliance while retaining robustness and localized controllability. Examples include integrating variable-stiffness metamaterials into the links [yitengma] or adopting compliant flexure-based joints instead of rigid pin joints. Importantly, we do not view uncontrolled compliance as a desirable goal, since it can lead to buckling, oscillation, and reduced payload capacity. Rather, the focus is on introducing compliance in a structured and tunable manner, so that the arm can selectively soften when adaptability is needed and stiffen when precision and force transmission are required.
In future work, we aim to address these limitations and extend UMArm’s control capabilities. We plan to work on the control aspect of the UMArm, which is particularly challenging due to the high degrees of freedom and the nonlinear behavior of soft actuators. Nonetheless, UMArm shows great potential: with a well-developed dynamical model and advanced motion planning, it could exploit its variable stiffness and compliance to achieve highly efficient, dynamic motions that are difficult to realize with conventional rigid robotic arms.
@@ VII Conclusion
This work introduces the UMArm, a pneumatically actuated rigid-soft hybrid arm that overcomes key limitations in soft robotic arms by integrating 24 independently controlled valve-embedded McKibben actuators. Our design achieves high precision (end-effector steady-state RMSE 0.85 mm), strong force output (53.3N while pulling at fully bend configuration), and rapid actuation while remaining lightweight (1.15 kg). Unlike traditional pneumatic arms requiring extensive external hardware, UMArm operates untethered using a compact, onboard pneumatic system, significantly enhancing portability. The modular architecture enables scalability and reconfiguration. The UMArm also demonstrates versatility. It can transform into an untethered inchworm robot that can move and interact. By leveraging antagonistic stiffening, the arm achieves directional compliance control. It can be tuned soft in one direction for safe human interaction, and rigid in the other direction for better force transmission. These advancements position UMArm as a robust solution for precise, wearable, and multimodal robotic applications. In the future, we would like to explore various control strategies that exploit the hardware advantages of the UMArm, striving for the goal of deploying high-performance, structurally compliant arms into the unstructured world.
VIII Acknowledgment
This research is funded by the University of Michigan, Mechanical Engineering Department.